Find the present value of an annuity with annual payments of $1,000.00 compounded at the end of each year for 9 years after being deferred for 4 years, if money is worth 6.5%, compounded annual.

The present value would be $_______
Answer #1

PVA = Annuity x [{1 - (1 + r)-n} / r]

= $1,000 x [{1 - 1.065-9} / 0.065] = $1,000 x [0.4326 / 0.065] = $1,000 x 6.6561 = $6,656.10

PV = PVA / (1 + r)^n = $6,656.10 / 1.065^4 = $6,656.10 / 1.2865 = $5,173.94